# Strong-to-Weak Spontaneous Symmetry Breaking in Mixed Quantum States

### v2

@@ Abstract
Symmetry in mixed quantum states can manifest in two distinct forms: strong symmetry, where each individual pure state in the quantum ensemble is symmetric with the same charge, and weak symmetry, which applies only to the entire ensemble. This paper explores a novel type of spontaneous symmetry breaking (SSB) where a strong symmetry is broken to a weak one. While the SSB of a weak symmetry is measured by the long-ranged two-point correlation function $\mathrm{Tr}(O_{x}O^{\dagger}_{y}\rho)$, the strong-to-weak SSB (SW-SSB) is measured by the fidelity $F(\rho,O_{x}O^{\dagger}_{y}\rho O_{y}O^{\dagger}_{x})$, dubbed the fidelity correlator. We prove that SW-SSB is a universal property of mixed-state quantum phases, in the sense that the phenomenon of SW-SSB is robust against symmetric low-depth local quantum channels. We also show that the symmetry breaking is “spontaneous” in the sense that the effect of a local symmetry-breaking measurement cannot be recovered locally. We argue that a thermal state at a nonzero temperature in the canonical ensemble (with fixed symmetry charge) should have spontaneously broken strong symmetry. Additionally, we study non-thermal scenarios where decoherence induces SW-SSB, leading to phase transitions described by classical statistical models with bond randomness. In particular, the SW-SSB transition of a decohered Ising model can be viewed as the “ungauged” version of the celebrated toric code decodability transition. We confirm that, in the decohered Ising model, the SW-SSB transition defined by the fidelity correlator is the only physical transition in terms of channel recoverability. We also comment on other (inequivalent) definitions of SW-SSB, through correlation functions with higher Rényi indices.
@@ II.1 Fidelity correlator and stability theorem
Again, if $\rho=\ket{\psi}\!\bra{\psi}$ |, then $\braket{O}_{F}=\left|\braket{\psi}{O}{\psi}\right|$ |, the modulus of the expectation value. But for a generic mixed state, the fidelity average is very different from the usual expectation value defined as $\braket{O}=\operatorname{Tr}[\rho O]$. If $O$ is unitary, then the above expression is the actual fidelity between two different density matrices $\rho$ and $\sigma:=O\rho O^{\dagger}$. If $O$ is charged under a strong symmetry of $\rho$, then $\langle O\rangle_{F}=0$.
@@
The proof of Theorem [1] can be proceeded till Eq. (11). Now the evolved operator $O^{\prime}\equiv UOU^{\dagger}$ acts nontrivially on two regions $X,Y$ (around the points $x,y$) with size $|X|,|Y|\sim D^{d}$, where $d$ is the spatial dimension. We can now decompose the operator $O^{\prime}$ as
@@
$D(\rho,\sigma)=\frac{1}{2}\operatorname{Tr}\|\rho-\sigma\|_{1}.$ (20)
@@ II.3 Local indistinguishability and detectability
here we suppose $O(x)$ is a unitary charged operator to ensure that $\tilde{\rho}$ is a legitimate density matrix, for simplicity [^2] The unitarity of $O(x)$ can be relaxed as long as we normalize Eq. (40) by $\operatorname{tr}\tilde{\rho}$. The subsequent arguments can be slightly modified to obtain essentially the same conclusion.. The presence of strong symmetry implies that $\rho$ and $\tilde{\rho}$ are orthogonal in the fidelity sense: $F(\rho,\tilde{\rho})=0$ because they carry different symmetry charges.
For any $k$-point observable (Eq. (38)), we have the following simple relation:
@@
$=$ $O(k/L^{d}).$
This means that any finite-point observable cannot distinguish $\tilde{\rho}$ from $\rho$. To distinguish the two states, the observable $M$ needs to act nontrivially on the entire system – the simplest such observable is the symmetry operator $g$.
@@
The state explicitly breaks the strong symmetry down to a weak symmetry. However, in general, this symmetry breaking may not be detectable using a local operator.
The symmetry breaking does become detectable using the fidelity average of a local charged operator if the original state $\rho$ has SW-SSB:
$F\left(\rho_{+},\frac{1}{L^{d}}\sum\limits_{y}O(y)\rho_{+}O^{\dagger}(y)\right)$
$\geq\frac{1}{L^{d}}\sum\limits_{y}F\left(\rho_{+},O(y)\rho_{+}O^{\dagger}(y)\right)$ (43)
$\geq\frac{1}{L^{2d}}\sum\limits_{x,y}F\left(\rho,O^{\dagger}(x)O(y)\rho O(x)O^{\dagger}(y)\right)\sim O(1).$
The inequalities above came from the joint concavity of fidelity:
It is straightforward to show that the above result on local detectability of explicit strong symmetry breaking for $\rho_{+}$ is also valid in the Rényi-$2$ sense.
@@ III.2 $\mathbb{Z}_{2}$ SW-SSB in the quantum Ising model
To detect the possible SW-SSB and identify the universality class of the symmetry-breaking transition, we consider the fidelity correlator of $Z$, which is charged under the strong $\mathbb{Z}_{2}$ symmetry. Then we utilize the replica trick to calculate the fidelity: the replicated fidelity is defined as
$F_{m,n}(x,y)=\operatorname{Tr}[(\rho^{m}\sigma\rho^{m})^{n}],$ (56)
here we utilized the following formula:
$\operatorname{Tr}(\rho^{t})\propto$ (62)
@@
Moreover, it is worth noting that the replicated fidelity in Eq. (56) also provides a naturally replicated version of the sandwiched Rényi relative entropy defined in Eq. (23), which can be recovered by taking the limit $m\to\frac{1-\alpha}{2\alpha}$ and $n\to\alpha$. Consequently, in the replica limit $t\to 1$, the partition function in Eq. (62) should also capture potential SW-SSB defined in terms of the sandwiched Rényi relative entropy. This offers supporting evidence for our earlier conjecture in Sec. II.1 that SW-SSB defined using distinct distinguishability measures – namely, quantum relative entropy and the fidelity correlator – ought to be equivalent.
@@
where $p_{n}\geq 0$ and $\sum_{n}p_{n}=1$. To determine the universality of SW-SSB transition, we again study the replicated partition function $\operatorname{Tr}\rho^{t}$. Let us first consider the loop (domain wall) basis, in which it is easier to identify the symmetry and the observable. The initial pure state can be represented in the domain wall basis as $\rho_{0}\propto\sum_{D,D^{\prime}}|D\rangle\langle D^{\prime}|$, where the summation is taken over all possible $\mathbb{Z}_{2}$ domain wall configurations. The decohered density matrix is
@@
$=\prod_{e}\beta^{\sum_{k}\frac{1}{2}\big{|}s_{e}^{(k)}-s_{e}^{(k+1)}\big{|}},$ (68)
exactly the same as the case of the Pauli channel (i.e. $\theta=\pi/2$), just with a different value of “tension”. This is true for each $t\geq 2$, and if we analytically continue to $t\rightarrow 1$, it makes sense to conclude that the partition function in the limit $t\rightarrow 1$ is still an RBIM. The model has a different value of $p_{c}$, but shows the same universality class of the SW-SSB transition.
@@ IV.2 Spontaneity and local irrecoverability
$D(\rho,\tilde{\rho})$$\geq$ $D(\rho,O_{j}^{\dagger}\rho O_{j})-D(\tilde{\rho},O_{j}^{\dagger}\rho O_{j})$
@@
where $\|O_{j}\|\sim O(1)$ is the operator norm of $O_{j}$. Putting all these relations back into Eq. (LABEL:eq:recovertracedistance), we obtain
@@
In this language, what we have shown in Thm. 4 is that the SSB of strong symmetry (either to weak or to trivial) is strongly spontaneous. Let us consider a pedagogical example of SW-SSB: $\rho_{0}=\mathbbm{1}+X$ as the SW-SSB of $\mathbb{Z}_{2}$ Ising symmetry (Sec. II.1). We perform a strong $Z$-measurement on a single qubit and recognize that $\rho_{0}$ will immediately collapse to the maximally mixed state $\mathbbm{1}$ that no longer has the strong Ising symmetry. Heuristically, the only information encoded in the SW-SSB density matrix is the global charge sector fixing, i.e., $\rho_{0}$ is the maximally mixed state within the even charge sector of the Ising symmetry. Once we perform a strong $Z$-measurement, the information on global charge sector fixing is completely destroyed, and we can never restore this information with any local quantum channel.
@@ V Summary and Outlook
    Strong-to-weak SSB of higher-form symmetry: A topologically ordered state can be understood as the spontaneous symmetry breaking of some higher-form symmetry Gaiotto et al. (2015); Wen (2019). Then what about the SW-SSB of higher-form symmetry? This question will be addressed in a forthcoming work Zhang et al. .
@@
    Ground state SSB comes with many dynamical features, such as Anderson tower and Goldstone modes (if the symmetry is continuous). It is natural to ask whether SW-SSB has similar dynamical consequences, for example in Lindbladian dynamics. Two prior works Ogunnaike et al. (2023); Ma and Turzillo (2024) addressed this issue partially, but a general picture is still lacking at this point.
@@ Acknowledgements.
We thank Yimu Bao, Tarun Grover, Timothy Hsieh, Andy Lucas, Zhu-Xi Luo, Max Metlitski, Shengqi Sang, Brian Swingle, Ruben Verresen, Xiao-Gang Wen, Yichen Xu, Carolyn Zhang, Zhehao Zhang, and Yijian Zou for inspiring discussions. RM is especially grateful to Shengqi Sang for many discussions and for explaining Ref. Sang and Hsieh (2024). JHZ and MC are grateful to Hong-Hao Tu and Wei Tang for discussions and collaborations on a related project. RM was supported in part by Simons Investigator Award number 990660 and by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (No. 651440). JHZ and ZB acknowledge support from NSF under award number DMR-2339319. RM, JHZ, ZB, and CW are supported in part by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. LAL acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute (LAL and CW) is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. MC acknowledges support from NSF under award number DMR-1846109. Note added – We would like to draw the reader’s attention to a related work Sala et al. on SW-SSB to appear in the same arXiv listing.

### v3

@@ Abstract
Symmetry in mixed quantum states can manifest in two distinct forms: strong symmetry, where each individual pure state in the quantum ensemble is symmetric with the same charge, and weak symmetry, which applies only to the entire ensemble. This paper explores a novel type of spontaneous symmetry breaking (SSB) where a strong symmetry is broken to a weak one. While the SSB of a weak symmetry is measured by the long-ranged two-point correlation function, the strong-to-weak SSB (SW-SSB) is measured by the fidelity correlator. We prove that SW-SSB is a universal property of mixed-state quantum phases, in the sense that the phenomenon of SW-SSB is robust against symmetric low-depth local quantum channels. We also show that the symmetry breaking is “spontaneous” in the sense that the effect of a local symmetry-breaking measurement cannot be recovered locally. We argue that a thermal state at a nonzero temperature in the canonical ensemble (with fixed symmetry charge) should have spontaneously broken strong symmetry. Additionally, we study non-thermal scenarios where decoherence induces SW-SSB, leading to phase transitions described by classical statistical models with bond randomness. In particular, the SW-SSB transition of a decohered Ising model can be viewed as the “ungauged” version of the celebrated toric code decodability transition. We confirm that, in the decohered Ising model, the SW-SSB transition defined by the fidelity correlator is the only physical transition in terms of channel recoverability. We also comment on other (inequivalent) definitions of SW-SSB, through correlation functions with higher Rényi indices.
@@ II.1 Fidelity correlator and stability theorem
Again, if $\rho=\ket{\psi}\!\bra{\psi}$ |, then $\braket{O}_{F}=\left|\braket{\psi}{O}{\psi}\right|$ |, the modulus of the expectation value. But for a generic mixed state, the fidelity average is very different from the usual expectation value defined as $\braket{O}=\operatorname{Tr}[\rho O]$. If $O$ is unitary, then the above expression is the actual fidelity between two different density matrices $\rho$ and $\sigma:=O\rho O^{\dagger}$. If $O$ is charged under a strong symmetry of $\rho$, then $\langle O\rangle_{F}=0$. Notice that
$\langle O\rangle_{F}$ $=$ $\operatorname{Tr}\sqrt{\sqrt{\rho}O\rho O^{\dagger}\sqrt{\rho}}=\|\sqrt{\rho}O\sqrt{\rho}\|_{1}$ (5)
$>$ $|\operatorname{Tr}\sqrt{\rho}O\sqrt{\rho}|=|\langle O\rangle|,$
which implies that if $\langle O\rangle\neq 0$, then we must have $\langle O\rangle_{F}>0$.
@@
We follow the same steps as the proof of Theorem 1 till Eq. (12). Now the evolved operator $O^{\prime}\equiv UOU^{\dagger}$ acts nontrivially on two regions $X,Y$ (around the points $x,y$) with size $|X|,|Y|\sim D^{d}$, where $d$ is the spatial dimension. We can now decompose the operator $O^{\prime}$ as
@@
$D(\rho,\sigma)=\frac{1}{2}\|\rho-\sigma\|_{1}.$ (21)
@@ II.2 Rényi-2 correlator
  3. 3.
    Experimentally measuring the fidelity correlator is extremely hard in the sense that it generally calls for the full quantum state tomography (QST) of the density matrix, whose complexity is exponential with respect to the system size. Therefore, the other significance of introducing the Rényi-2 correlator is the experimental accessibility: the Rényi-2 correlator can be experimentally measured without performing QST. Recently, the proposed classical shadow tomography Huang et al. (2020); Elben et al. (2022); Hu et al. (2023) significantly reduces the experimental complexity of measuring nonlinear quantum information-theoretic quantities, especially in open quantum systems. In particular, as indicated in Eq. (30), both the numerator and denominator of the Rényi-2 correlator are polynomials of the density matrix $\rho$, which emphasizes that both can be efficiently measured by classical shadow tomography.
@@ II.3 Local indistinguishability and detectability
here we suppose $O(x)$ is a unitary charged operator to ensure that $\tilde{\rho}$ is a legitimate density matrix, for simplicity. The presence of strong symmetry implies that $\rho$ and $\tilde{\rho}$ are orthogonal in the fidelity sense: $F(\rho,\tilde{\rho})=0$ because they carry different symmetry charges.
For any $k$-point observable (Eq. (39)), we have the following simple relations:
@@
$=$ $O(k/L^{d}),$
and
$F\left(\rho,M_{k}\rho M_{k}^{\dagger}\right)-F\left(\tilde{\rho},M_{k}\tilde{\rho}M_{k}^{\dagger}\right)\leq F\left(\rho,M_{k}\rho M_{k}^{\dagger}\right)-\frac{1}{L^{d}}\sum_{x}F\left(O_{x}\rho O_{x}^{\dagger},M_{k}O_{x}\rho O_{x}^{\dagger}M_{k}^{\dagger}\right)$
$=$ $F\left(\rho,M_{k}\rho M_{k}^{\dagger}\right)-\frac{1}{L^{d}}\sum_{x}F\left(O_{x}\rho O_{x}^{\dagger},\left([M_{k},O_{x}]+O_{x}M_{k}\right)\rho\left(M_{k}^{\dagger}O_{x}^{\dagger}+\left[O_{x}^{\dagger},M_{k}^{\dagger}\right]\right)\right)$
$=$ $\frac{1}{L^{d}}\sum_{x\in k}F\left(\rho,M_{k}\rho M_{k}^{\dagger}\right)-F\left(O_{x}\rho O_{x}^{\dagger},\left([M_{k},O_{x}]+O_{x}M_{k}\right)\rho\left(M_{k}^{\dagger}O_{x}^{\dagger}+\left[O_{x}^{\dagger},M_{k}^{\dagger}\right]\right)\right)\sim O\left(\frac{k}{L^{d}}\right),$ (43)
where $x\in k$ means that the site $x$ locates inside the support of $M_{k}$. The inequalities above came from the joint concavity of fidelity:
We conclude that any finite-point observable cannot distinguish $\tilde{\rho}$ from $\rho$. To distinguish the two states, the observable $M$ needs to act nontrivially on the entire system – the simplest such observable is the symmetry operator $g$.
@@
The state explicitly breaks the strong symmetry down to a weak symmetry.
The symmetry breaking does become locally detectable using the fidelity average of a local charged operator if the original state $\rho$ has SW-SSB:
$F\left(\rho_{+},O^{\dagger}(y)\rho_{+}O(y)\right)$
$\geq\frac{1}{L^{d}}\sum_{x}F\left(\rho,O(x)O^{\dagger}(y)\rho O(y)O^{\dagger}(x)\right)\sim O(1),$ (46)
where the inequality is ensured by the joint concavity of the fidelity (44). It is straightforward to show that the above result on local detectability of explicit strong symmetry breaking for $\rho_{+}$ is also valid in the Rényi-$2$ sense.
@@ III.2 $\mathbb{Z}_{2}$ SW-SSB in the quantum Ising model
To detect the possible SW-SSB and identify the universality class of the symmetry-breaking transition, we consider the fidelity correlator of $Z$, which is charged under the strong $\mathbb{Z}_{2}$ symmetry.
$\rho=\sum_{\braket{X}}Z(X)\ket{X}\bra{X}.$ | . (58)
$\sigma$ $=Z_{i}Z_{j}\rho Z_{i}Z_{j}$
$=\sum_{X}Z(X\cdot\delta X)\ket{X}\bra{X}.$ | . (62)
$F_{ij}$ $=\operatorname{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$ (63)
$=\sum_{X}\sqrt{Z(X)Z(X\cdot\delta X)}$
$=\sum_{X}Z(X)\left(\frac{Z(X\cdot\delta X)}{Z(X)}\right)^{1/2}$
$=\sum_{X}Z(X)e^{-\frac{1}{2}\Delta F_{\mathrm{d}w}}$
Here $\Delta F_{\rm dw}=-\ln\frac{Z(X\cdot\delta X)}{Z(X)}$ is the free energy cost of inserting a domain wall of length $\sim|i-j|$. $\braket{\cdots}_{\rm RBIM}$ stands for disorder average in the random-bond Ising model along the Nishimori line Dennis et al. (2002). For $p>p_{c}$, the RBIM is in the paramagnetic phase (i.e. high temperature), so the disorder-averaged free energy cost for a domain wall approaches a constant. Thus we have shown that the fidelity correlator is long-range ordered for $p>p_{c}$. It also suggests that $F_{ij}$ decays exponentially for $p<p_{c}$, since the free energy cost diverges linearly with the distance in the ferromagnetic (low-temperature) phase.
We note that the same results follow for the other definition of SSB in terms of sandwiched Rényi entropy. offering supporting evidence for our earlier conjecture in Sec. II.1 that SW-SSB defined using distinct distinguishability measures – namely, quantum relative entropy and the fidelity correlator – ought to be equivalent.
@@
where $p_{n}\geq 0$ and $\sum_{n}p_{n}=1$.
For these non-Pauli channels, the density matrix is no longer diagonal in the $X$ basis. In the following, we will argue that nevertheless the Ising SW-SSB transition is still described by the RBIM universality class, the same as the Pauli case. We will utilize the replica trick to calculate the fidelity, which is not rigorous but still suggestive. In Appendix B we perform the calculations in the Pauli case $\theta=\pi/2$, and reproduce the same result as the exact calculation above, providing evidence for the validity of the replica trick.
We define the replicated fidelity as
$F_{m,n}(x,y)=\operatorname{Tr}[(\rho^{m}\sigma\rho^{m})^{n}],$ (67)
and the standard definition of fidelity (7) is recovered by taking the limit $m,n\rightarrow 1/2$. Thus to compute $F_{m,n}(x,y)$, we need to study the replicated partition function $\operatorname{Tr}\rho^{t}$. It is not difficult to see that $t=(2m+1)n$, so the limit $m,n\rightarrow 1/2$ corresponds to $t\rightarrow 1$.
Let us consider the domain wall basis (i.e. eigenbasis of $Z$), in which it is easier to compute the matrix elements of $\rho$. The initial pure state can be represented in the domain wall basis as $\rho_{0}\propto\sum_{D,D^{\prime}}|D\rangle\langle D^{\prime}|$, where the summation is taken over all possible $\mathbb{Z}_{2}$ domain wall configurations. The decohered density matrix is
@@
$=\prod_{e}\beta^{\sum_{k}\frac{1}{2}\big{|}s_{e}^{(k)}-s_{e}^{(k+1)}\big{|}}.$ (71)
Crucially, the replicated partition function takes the same form regardless of the value of $\theta$, just with different values of the “tension” $\beta$. This is true for each $t\geq 2$, and if we analytically continue to $t\rightarrow 1$, it makes sense to conclude that the partition function in the limit $t\rightarrow 1$ should have the same universal behavior as the $\theta=\pi/2$ case. Thus if the replica limit exists, the SW-SSB transition in the non-Pauli model defined in Eq. (65) is described by the same RIBM universality class.
@@ IV.2 Spontaneity and local irrecoverability
$D(\rho,\tilde{\rho})\geq D(\rho,O_{j}^{\dagger}\rho O_{j})-D(\tilde{\rho},O_{j}^{\dagger}\rho O_{j})$ (95)
@@
where $\|O_{j}\|\sim O(1)$ is the operator norm of $O_{j}$. Putting all these relations back into Eq. (95), we obtain
@@
In this language, what we have shown in Thm. 4 is that the SSB of strong symmetry (either to weak or to trivial) is strongly spontaneous. Let us consider a pedagogical example of SW-SSB: $\rho_{0}\propto\mathbbm{1}+X$ as the SW-SSB of $\mathbb{Z}_{2}$ Ising symmetry (Sec. II.1). We perform a strong $Z$-measurement on a single qubit and recognize that $\rho_{0}$ will immediately collapse to the maximally mixed state $\mathbbm{1}$ that no longer has the strong Ising symmetry. Heuristically, the only information encoded in the SW-SSB density matrix is the global charge sector fixing, i.e., $\rho_{0}$ is the maximally mixed state within the even charge sector of the Ising symmetry. Once we perform a strong $Z$-measurement, the information on global charge sector fixing is completely destroyed, and we can never restore this information with any local quantum channel.
@@
For pure states, the spontaneity of SSB is quantitatively described by a divergent dynamical susceptibility. For example, the magnetic susceptibility of the GHZ state is divergent with respect to the external Zeeman field along the $Z$ direction. Following the initial version of our manuscript, in Ref. Zhang et al. (2024a) the authors proposed a fidelity susceptibility which diverges in the SW-SSB phase and remains finite in the symmetric phase.
@@ V Summary and Outlook
    Strong-to-weak SSB of higher-form symmetry: A topologically ordered state can be understood as the spontaneous symmetry breaking of some higher-form symmetry Gaiotto et al. (2015); Wen (2019). Then what about the SW-SSB of higher-form symmetry? This question will be addressed in a forthcoming work Zhang et al. (2024b).
@@
    Ground state SSB comes with many dynamical features, such as Anderson tower and Goldstone modes (if the symmetry is continuous). It is natural to ask whether SW-SSB has similar dynamical consequences, for example in Lindbladian dynamics. Two prior works Ogunnaike et al. (2023); Ma and Turzillo (2024) addressed this issue partially, but a general picture is still lacking at this point. In particular, a follow-up work Huang et al. (2024) showed that the Goldstone mode of U(1) SW-SSB can be identified as the diffusive hydrodynamic mode, using Schwinger-Keldysh effective field theory.
@@ Acknowledgements.
We thank Yimu Bao, Tarun Grover, Timothy Hsieh, Andy Lucas, Zhu-Xi Luo, Max Metlitski, Shengqi Sang, Brian Swingle, Ruben Verresen, Xiao-Gang Wen, Yichen Xu, Carolyn Zhang, Zhehao Zhang, and Yijian Zou for inspiring discussions. RM is especially grateful to Shengqi Sang for many discussions and for explaining Ref. Sang and Hsieh (2024). JHZ and MC are grateful to Hong-Hao Tu and Wei Tang for discussions and collaborations on a related project. RM was supported in part by Simons Investigator Award number 990660 and by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (No. 651440). JHZ and ZB acknowledge support from NSF under award number DMR-2339319. RM, JHZ, ZB, and CW are supported in part by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. LAL acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute (LAL and CW) is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. MC acknowledges support from NSF under award number DMR-1846109. Note added – We would like to draw the reader’s attention to a related work Sala et al. (2024) on SW-SSB to appear in the same arXiv listing.